# Sound-VECaps: IMPROVING AUDIO GENERATION WITH VISUAL ENHANCED CAPTIONS

### v3

@@ 5 Conclusion
Experiments show that AudioLDM models trained on Sound-VECaps achieve SoTA performance and outperform baseline models. In addition, our systems further improve by a large margin when taking more detailed captions as prompts, reaching a FAD score of . In addition, Sound-VECaps demonstrate substantial improvements in audio retrieval and temporal feature identification. Nevertheless, the results between AudioCaps and AudioCaps-Enhanced testing sets highlight the limitations of previous benchmarks that rely on simple prompts and emphasize the value and potential of the more detailed prompts in advancing the performance of audio-language models. For both the Sound-VECaps and AudioCaps-Enhanced testing set, we developed audio-only versions and full versions with captions that include and exclude visual-only content for different purposes and tasks. Although the improvements from the dataset are confined to specific audio-language tasks, we hope these datasets can offer advancements in audio generation systems and support the community in making future progress in audio-language tasks.
6 ACKNOWLEDGMENT
@@
This research was partly supported by a research scholarship from the China Scholarship Council (CSC), an internship from ByteDnace, and supported by the Engineering and Physical Sciences Research Council (EPSRC) under Grant EP/T019751/1 “AI for Sound”.
@@ A.1 LLMs Prompts
A.2 Caption Demos
We present the comparison of the captions from Sound-VECaps and other baseline datasets. As shown in Table 6, for each audio sample, we compare the caption from the AudioSet label, Wavcaps, Enclap, Auto-ACD and two versions of the proposed Sound-VECaps.
Nevertheless, we also present a sample of the proposed AudioCaps-Enhanced testing dataset in Table 7

### v4

@@ V Conclusion
We present Sound-VECaps, a large-scale dataset comprising M audio clips with captions augmented by video data, to address the challenge of prompt following in audio generation systems. Experiments show that systems trained on Sound-VECaps achieve SoTA performance and outperform baseline models. In addition, a new benchmark using improved captions is proposed to evaluate audio-language systems on complex and extended prompts. Our systems are further improved by a large margin when taking more detailed captions as prompts, reaching a FAD score of . Nevertheless, we demonstrated that using Sound-VECaps can offer substantial improvements in audio-language and temporal feature retrieval. The results of the proposed AudioCaps-Enhanced testing sets highlight the limitations of previous benchmarks and emphasize the potential of better prompts in advancing the performance of audio-language models. Overall, we develop two versions of the proposed datasets with captions that include and exclude visual-only content for different purposes and tasks and hope these datasets will generate more profound impacts on audio-language learning.
VI ACKNOWLEDGMENT
@@
This research was partly supported by a research scholarship from the China Scholarship Council (CSC), an internship from ByteDance, and supported by the Engineering and Physical Sciences Research Council (EPSRC) under Grant EP/T019751/1 “AI for Sound”. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) license to any Author Accepted Manuscript version arising.
@@ A-A LLMs Prompts
A-B Caption Demos
We present the comparison of the captions from Sound-VECaps and other baseline datasets. As shown in Table VI, for each audio sample, we compare the caption from the AudioSet label, Wavcaps, Enclap, Auto-ACD and two versions of the proposed Sound-VECaps. A sample of the proposed AudioCaps-Enhanced testing dataset is presented in Table VII.